A golf ball has a mass of .046 grams and when struck by a driver reaches a speed of 42 m/s. The club is in contact with the ball for the time of .50 millisecond. If the momentum of the ball just after the impact is 1.9 m/s, what is the average force of the club on the ball?
1
Expert's answer
2012-11-22T09:45:26-0500
The average force F=dp/dt=((V1-V2)*m)/t=((42m/s-1.9 m/s)*0.000046 kg)/0.0005 s=3.6892 N